OpenStudy (anonymous):

Write an equation of the line that passes through the given point and is (a) parallel to the given line and (b) perpendicular to the given line. (2,-3), y= -2x -3

OpenStudy (mertsj):

y-(-3)=-2(x-2) is the parallel line y-(-2)=1/2(x-2) is the perpendicular